# Precise quantum-geometric electronic properties from first principles

### v1

@@ 6 Conclusions
The numerically precise calculation of the basic quantum-geometric properties of crystals, namely Berry curvature, quantum metric, orbital magnetic moment and effective mass, was successfully implemented in a first-principles pseudopotential plane-wave code. The adopted procedure is as follows. For a given effective potential and wavevector ${\bf k}$, the first derivatives of the cell-periodic Bloch wavefunctions $u_{dn}({\bf k})$ with respect to wavevector ${\bf k}$ are obtained from perturbation theory. This is achieved by solving iteratively a Sternheimer equation, taking into account energy-level degeneracies and using a stable algorithm. Once those wavefunction derivatives are obtained, the four quantum-geometric quantities are readily determined as the real and imaginary parts of two complex objects: the quantum-geometric tensor $T^{\alpha\beta}_{dd^{\prime}}$ of Eq. (22) (quantum metric and the Berry curvature), and the closely-related mass-moment tensor $\Gamma^{\alpha\beta}_{dd^{\prime}}$ of Eq. (28) (inverse effective mass and orbital moment).
@@
In both Si and GaAs, the top of the valence band has nontrivial degeneracies. As a result, the $T$ and $\Gamma$ tensors have a total of four indices: two for the spatial dimensions ($\alpha\beta$), and two for the level degeneracy ($dd^{\prime}$). The calculation of the direction-dependent transport-equivalent effective masses illustrates how to extract physical properties from those complex objects, taking into account the subtleties of perturbation theory for degenerate states.
Calculations of quantum-geometric quantities for gapped graphene were performed by adding a bespoke artificial potential to first-principles graphene. As the two bands near the opened gap are reasonably decoupled from other bands, in particular once mirror-symmetry selection rules are accounted for, the first-principles results can be compared to a well-known two-band low-energy model. With appropriate parameters, that model closely reproduced the calculated DFT values, apart from some deviations that were discussed. Such detailed comparison between low-energy models and direct ab initio calculations has not, to our knowledge, been reported in the literature.
Trigonal tellurium, one of the simplest crystals with interesting quantum geometric properties, was chosen as the final example. The properties were calculated in the neighborhood of the valence and conduction band edges. For the “camel back” region of the valence-band maximum, the calculated orbital magnetic moment was similar to the results of a previous calculation. In the low-lying conduction bands, the orbital moment near the H point was found to have the opposite sign from that of the upper valence band. This could have interesting implications for nonlinear magneto-transport properties such as electrical magnetochiral anisotropy [59].
@@ Acknowledgments
Work by J. L. M. and C. L. R. was supported by grant EXPL-FIS-MAC-1334-2021 from the Portuguese Science and Technology Foundation (FCT). They wish to acknowledge FCT for funding the Research Unit INESC MN (UID/05367/2020) through Plurianual, Base and Programatico financing. Work by I. S. was supported by Grant No. PID2021-129035NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by ERDF/EU. The authors thank C. J. Augusto of Quantum Semiconductor LLC for useful discussions throughout this work, and for a careful reading of the manuscript.

### v2

@@ 6 Conclusions
The numerically precise calculation of the basic quantum-geometric properties of crystals, namely Berry curvature, quantum metric, orbital magnetic moment and effective mass, was successfully implemented in a first-principles pseudopotential plane-wave code. The adopted procedure is as follows. For a given effective potential and wavevector $\bf k$}, the first derivatives of the cell-periodic Bloch wavefunctions $u_{dn{\bf k}}({\bf r})$ with respect to wavevector $\bf k$} are obtained from perturbation theory. This is achieved by solving iteratively a Sternheimer equation, taking into account energy-level degeneracies and using a stable algorithm. Once those wavefunction derivatives are obtained, the four quantum-geometric quantities are readily determined as the real and imaginary parts of two complex objects: the quantum-geometric tensor $T^{\alpha\beta}_{dd^{\prime}}$ of Eq. (22) (quantum metric and the Berry curvature), and the closely-related mass-moment tensor $\Gamma^{\alpha\beta}_{dd^{\prime}}$ of Eq. (28) (interband inverse effective mass and orbital magnetic moment). The intraband part contribution to the inverse effective mass given by matrix elements of the operator $H^{\alpha\beta}_{\bf k}$, is then included according to Eq. (30).
@@
In both Si and GaAs, the top of the valence band has nontrivial degeneracies. As a result, the $T$ and $\Gamma$ tensors have a total of four indices: two for the spatial dimensions ($\alpha\beta$), and two for the level degeneracy ($dd^{\prime}$). The calculation of the direction-dependent transport equivalent effective masses illustrates how to extract physical properties from those complex objects, taking into account the subtleties of perturbation theory for degenerate states.
Calculations of quantum-geometric quantities for gapped graphene were performed by adding a bespoke artificial potential to first-principles graphene. As the two bands near the opened gap are reasonably decoupled from other bands, in particular once mirror-symmetry selection rules are accounted for, the first-principles results can be compared to a well-known two-band low-energy model. With appropriate parameters, that model closely reproduced the calculated DFT values, apart from some small deviations that were discussed. Such detailed comparison between low-energy models and direct ab initio calculations has not, to our knowledge, been reported in the literature.
Trigonal tellurium, one of the simplest crystals with interesting quantum-geometric properties, was chosen as the final example. The properties were calculated in the neighborhood of the valence-band and conduction-band edges. For the camel-back region of the valence-band maximum, the calculated orbital magnetic moment was similar to the results of a previous calculation. In the low-lying conduction bands, the orbital moment near the H point was found to have the opposite sign from that of the upper valence band. This could have interesting implications for nonlinear magneto-transport properties such as electrical magnetochiral anisotropy [60].
@@ Acknowledgments
Work by J. L. M. and C. L. R. was supported by grant EXPL-FIS-MAC-1334-2021 from the Portuguese Science and Technology Foundation (FCT). They wish to acknowledge FCT for funding the Research Unit INESC MN (UID/05367/2020) through Plurianual, Base and Programático financing. Work by I. S. was supported by Grant No. PID2021-129035NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by ERDF/EU. The authors thank C. J. Augusto of Quantum Semiconductor LLC for useful discussions throughout this work, and for a careful reading of the manuscript.